# Enhanced Phase Estimation via Photon-Added Two-Mode Squeezed States and Kerr Nonlinearity

### v1

@@ III Phase Sensitivity of Intensity Difference Detection
Figs. 4(a)-(c) depict linear phase shift sensitivity $\Delta\phi$ as a function of loss rate $l$, compared with SQL and HL. Despite photon loss, phase sensitivity can still surpass SQL, with photon addition improving system robustness. Figs. 5(a)-(c) further illustrate Kerr nonlinear phase shift sensitivity versus loss rate $l$. Comparison with Fig. 4 clearly shows that Kerr nonlinearity enhances phase sensitivity by an order of magnitude, enabling exceedance of the HL threshold even with loss, whereas linear phase shift only surpasses SQL within limited $l$ ranges. However, with fixed $r=0.9$ and $\alpha=1$, phase sensitivity for $m=n=0$ can exceed sub-HL, but not for $m=n=1,2$. This may arise from Kerr nonlinearity sensitivity to other parameters (e.g., phase shift). This situation indicates that without photon addition, phase sensitivity can exceed sub-HL within certain ranges, which is relatively favorable for Kerr nonlinearity.
@@ IV.1 QFI of KMZI in Ideal Case
Quantum Fisher information (QFI) quantifies the maximum information about phase shift $\phi$ that a quantum state can convey after passing through a phase shifter. In our scheme, employing linear and Kerr nonlinear phase shifters within the MZI leads to two corresponding QFI representations, denoted $F_{k}$ ($k=1,2$). QFI is independent of any specific detection scheme and provides an upper bound on classical Fisher information obtainable through $\phi$ estimation. The optimal error bound for phase sensitivity, known as the quantum Cramér-Rao bound (QCRB), is given by 44 ; 45 ; 46 :
@@
$F_{k}=4\left[\left\langle\psi_{\phi}^{\prime}|\psi_{\phi}^{\prime}\right\rangle-\left|\left\langle\psi_{\phi}^{\prime}|\psi_{\phi}\right\rangle\right|^{2}\right],$ (14)
where $\left|\psi_{\phi}\right\rangle=U\left(\phi,k\right)B_{1}\left|\psi\right\rangle_{\text{in}}$ represents the quantum state before BS2 in the lossless KMZI, and $\left|\psi_{\phi}^{\prime}\right\rangle=\partial\left|\psi_{\phi}\right\rangle/\partial\phi$. For both linear ($k=1$) and Kerr nonlinear ($k=2$) phase shifts, QFI can be further simplified as 40 :
$F_{k}=4\left\langle\Delta^{2}n_{b}^{k}\right\rangle=4\left[\left\langle\left(n_{a}^{k}\right)^{2}\right\rangle-\left\langle n_{a}^{k}\right\rangle^{2}\right],$ (15)
where $\left\langle\cdot\right\rangle=\left.\left\langle\Psi\right|\cdot\left|\Psi\right\rangle\right$., with $\left|\Psi\right\rangle=B_{1}\left|\text{in}\right\rangle$. Here, $n_{a}^{k}=\left(a^{\dagger}a\right)^{k}$ ($k=1,2$). To facilitate QFI computation, the normal ordering form of $n_{a}^{w}$ is obtained via the operator identity:
$n_{a}^{w}$ $=$ $\frac{\partial^{w}}{\partial x^{w}}\left.\exp\left[xa^{\dagger}a\right]\right|_{x=0}$ (16)
$=$ $\frac{\partial^{w}}{\partial x^{w}}\left.\colon\exp\left[\left(e^{x}-1\right)a^{\dagger}a\right]\colon\right|_{x=0}.$
@@
$F_{1}=4\left[D_{1}\left(2,0,2,0\right)+D_{1}\left(1,0,1,0\right)-D_{1}\left(1,0,1,0\right)^{2}\right],$ (17)
@@
$F_{2}$ $=$ $4\left[D_{1}\left(4,0,4,0\right)+6D_{1}\left(3,0,3,0\right)\right$. (18)
$+7D_{1}\left(2,0,2,0\right)+D_{1}\left(1,0,1,0\right)$
$\left.-\left(D_{1}\left(2,0,2,0\right)+D_{1}\left(1,0,1,0\right)\right)^{2}\right].$
Refer to captionRefer to captionRefer to captionRefer to caption
@@
As evident from Fig. 6, increasing photon addition order significantly enhances QFI. Furthermore, comparing Figs. 6(a)-(b) with Figs. 6(c)-(d) clearly shows that the $k=2$ case yields larger QFI values. This demonstrates that both photon addition operations and Kerr nonlinear phase shifts effectively improve QFI.
@@ IV.2.1 Effects of Photon Loss on QFI in Linear Phase Shift Case ($k=1$)
Practical measurements involve photon losses due to vacuum noise, affecting QFI. For the initial pure state $\left|\Psi_{S}\right\rangle$ ($\left|\Psi_{S}\right\rangle=B_{1}\left|\psi\right\rangle_{\text{in}}$) in a probe system $S$ with a lossy MZI, we introduce environment $E$ orthogonal states $\left|j_{E}\right\rangle$ and Kraus operators $\hat{\Pi}_{j}\left(\phi\right)$ to characterize $\left|\Psi_{S}\right\rangle$ evolution through the phase shifter with photon losses. Fig. 7 illustrates the photon loss within the MZI, providing a visual representation of the model structure under investigation. Following Escher et al.’s methods for calculating QFI in open quantum systems 49 , the quantum states $\left|\Psi_{S}\right\rangle$ and environment vacuum noise $\left|0_{E}\right\rangle$, after unitary evolution $U_{S+E}\left(\phi\right)$, can be expressed in the extended $S+E$ space:
$\left|\Psi_{S+E}\right\rangle$ $=$ $U_{S+E}\left(\phi\right)\left|\Psi_{S}\right\rangle\left|0_{E}\right\rangle$ (19)
$=$ $\sum_{j=0}^{\infty}\hat{\Pi}_{j}\left(\phi\right)\left|\Psi_{S}\right\rangle\left|j_{E}\right\rangle.$
Although photon losses transform $\left|\Psi_{S}\right\rangle$ into a mixed state, this approach allows treating $\left|\Psi_{S+E}\right\rangle$ in $S+E$ as a pure state. For the entire purified system, QFI under photon losses can be expressed as:
$F_{L_{1}}\leqslant C_{Q}\left[\left|\Psi_{S}\right\rangle,\hat{\Pi}_{j}\left(\phi\right)\right]=4\left[\left\langle\hat{H}_{1}\right\rangle_{S}-\left|\left\langle\hat{H}_{2}\right\rangle_{S}\right|^{2}\right],$ (20)
where the lower bound of $C_{Q}\left[\left|\Psi_{S}\right\rangle,\hat{\Pi}_{j}\left(\phi\right)\right]$ is shown to be the QFI for the reduced system 49 , and $\hat{H}_{1,2}$ are Hermitian operators defined as:
$\hat{H}_{1}$ $=$ $\sum\limits_{j=0}^{\infty}\frac{d\hat{\Pi}_{j}^{\dagger}\left(\phi\right)}{d\phi}\frac{d\hat{\Pi}_{j}\left(\phi\right)}{d\phi},$ (21)
$\hat{H}_{2}$ $=$ $i\sum\limits_{j=0}^{\infty}\frac{d\hat{\Pi}_{j}^{\dagger}\left(\phi\right)}{d\phi}\hat{\Pi}_{j}\left(\phi\right),$ (22)
with $\left\langle\cdot\right\rangle_{S}$ denoting expectation value $\left.\left\langle{}_{S}\Psi\right|\cdot\left|\Psi_{S}\right\rangle\right$., representing the average of $\hat{H}_{1,2}$. Thus, $F_{L_{1}}=C_{Q\min}$, and for linear phase shift ($k=1$), the Kraus operator is:
$\hat{\Pi}_{j}\left(\phi\right)=\sqrt{\frac{l^{j}}{j!}}e^{i\phi\left(a^{\dagger}a-\gamma j\right)}\left(1-l\right)^{\frac{a^{\dagger}a}{2}}a^{j},$ (23)
@@
$F_{L_{1}}=\frac{4F_{1}\left(1-l\right)\left\langle n_{a}\right\rangle}{lF_{1}+4\left(1-l\right)\left\langle n_{a}\right\rangle}.$ (24)
@@ IV.2.2 Effects of Photon Loss on QFI in Kerr Nonlinear Phase Shift Case ($k=2$)
$\hat{\Pi}_{j}\left(\phi\right)=\sqrt{\frac{l^{j}}{j!}}e^{i\phi\left[\left(a^{\dagger}a\right)^{2}-2\mu_{1}a^{\dagger}aj-\mu_{2}j^{2}\right]}\left(1-l\right)^{\frac{a^{\dagger}a}{2}}a^{j},$ (25)
@@
$F_{L_{2}}$ $\leqslant$ $C_{Q}\left[\left|\Psi_{S}\right\rangle,\hat{\Pi}_{j}\left(\phi\right)\right]$ (26)
$=$ $4\left[K_{1}^{2}\left\langle\Delta^{2}n_{a}^{2}\right\rangle_{S}-K_{2}\left\langle n_{a}^{3}\right\rangle_{S}\right$.
$+K_{3}\left\langle n_{a}^{2}\right\rangle_{S}-K_{4}\left\langle n_{a}\right\rangle_{S}$
$\left.-K_{5}\left\langle n_{a}^{2}\right\rangle_{S}\left\langle n_{a}\right\rangle_{S}-K_{6}\left\langle n_{a}\right\rangle_{S}^{2}\right],$
@@ V Conclusion
This study demonstrates that injecting PA-TMSCS into a KMZI significantly enhances phase measurement precision. By combining photon addition operations with Kerr nonlinear phase shift, we achieved measurement precision surpassing SQL and approaching sub-HL and SHL performance levels. Systematic analysis confirms that increasing photon addition order and input resource strength enhances both phase sensitivity and QFI. Crucially, the scheme maintains enhanced robustness against photon loss, providing a promising pathway toward practical quantum-enhanced metrology applications. The synergistic combination of photon addition operations and Kerr nonlinear phase shifting offers a powerful approach for achieving ultra-high precision measurements in practical quantum metrology systems.
Acknowledgements.
@@
This work is supported by the National Natural Science Foundation of China (Grants No. 12564049 and No. 12104195) and the Jiangxi Provincial Natural Science Foundation (Grants No. 20242BAB26009 and 20232BAB211033), Jiangxi Provincial Key Laboratory of Advanced Electronic Materials and Devices (Grant No. 2024SSY03011), as well as Jiangxi Civil-Military Integration Research Institute (Grant No. 2024JXRH0Y07).
@@
APPENDIX A : DERIVATION OF TRANSFORMATION RELATION FOR KERR NONLINEAR PHASE SHIFTER OPERATOR
@@
In the scheme proposed in this paper, the Kerr nonlinear phase shifter significantly enhances phase measurement accuracy. The associated calculations can be efficiently performed using the transformation relation of the equivalent operator of the Kerr nonlinear phase shifter, $U\left(\phi,2\right)=e^{i\phi\left(a^{\dagger}a\right)^{2}}$, as shown in our study. Therefore, this appendix provides a detailed derivation of the transformation relation $U^{\dagger}\left(\phi,2\right)bU\left(\phi,2\right)=e^{i\phi}e^{i\phi\left(2a^{\dagger}a\right)}a$ in Eq. (8). Here, we assume the operator corresponding to the transformation relation of $U\left(\phi,2\right)$ is $O$, i.e., $O=U^{\dagger}\left(\phi,2\right)bU\left(\phi,2\right)$. The matrix of $O$ in the Fock state space can be represented by $O_{n,n^{\prime}}=\left\langle n\right|O\left|n^{\prime}\right\rangle$, and its corresponding expansion is given by
$O=\sum_{n,n^{\prime}=0}^{\infty}\left|n\right\rangle\left\langle n^{\prime}\right|O_{n,n^{\prime}}.$ (A1)
@@
By utilizing $\left\langle n\right|b^{\dagger}b=\left\langle n\right|n$ and $\left\langle n\right|b=\left\langle n+1\right|\sqrt{n+1}$, the matrix element $O_{n,n^{\prime}}$ in the aforementioned equation can be further calculated as
$O_{n,n^{\prime}}=\sqrt{n+1}e^{i\phi\left(2n+1\right)}\delta_{n+1,n^{\prime}},$ (A2)
and by substituting it into Eq. (A1), the transformation relation of $U\left(\phi,2\right)$ can be deduced as follows
$O=$ $e^{i\phi}\sum_{n=0}^{\infty}\left|n\right\rangle\left\langle n+1\right|\sqrt{n+1}e^{i2n\phi}$
$=e^{i\phi}\sum_{n=0}^{\infty}\left|n\right\rangle\left\langle n\right|e^{i2n\phi}a=e^{i\phi}e^{i\phi\left(2a^{\dagger}a\right)}a,$ (A3)
@@
where we have used the completeness of the Fock state $\sum\limits_{n=0}^{\infty}\left|n\right\rangle\left\langle n\right|=1$.
@@
APPENDIX B : THE SPECIFIC DERIVATIONS AND EXPRESSIONS FOR $\left\langle I_{D}\right\rangle$ AND $\left\langle I_{D}^{2}\right\rangle$ UNDER $k=1$ CONDITION
@@
In order to obtain the mean values $\left\langle I_{D}\right\rangle$ and $\left\langle I_{D}^{2}\right\rangle$ in Eq. (10) in a simple way, the general formula is introduced i.e.,
$_{f}\left\langle 0\right|\left\langle in\right|B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi\left(a^{\dagger}a\right)}$
$\times\left(a^{\dagger m_{1}}b^{\dagger n_{1}}a^{m_{2}}b^{n_{2}}\right)e^{i\phi\left(a^{\dagger}a\right)}B_{f}B_{1}\left|in\right\rangle\left|0\right\rangle_{f}$
$=D_{1}\left(m_{1},n_{1},m_{2},n_{2}\right).$ (B1)
@@
The quantum state used in the above formula to calculate the average value is the state after passing through the phase shifter, and the output state still needs to pass through BS2. Thus using the transformation relation Eqs. (6) and (B1) one can obtain $\left\langle I_{D}\right\rangle$ and $\left\langle I_{D}^{2}\right\rangle$ respectively as
$\left\langle I_{D}\right\rangle=i\left[D_{1}\left(1,0,0,1\right)-D_{1}\left(0,1,1,0\right)\right],$ (B2)
@@
$\left\langle I_{D}^{2}\right\rangle$ $=D_{1}\left(0,1,0,1\right)+D_{1}\left(1,0,1,0\right)+2D_{1}\left(1,1,1,1\right)$
$-D_{1}\left(2,0,0,2\right)-D_{1}\left(0,2,2,0\right).$ (B3)
@@
$D_{1}\left(m_{1},n_{1},m_{2},n_{2}\right)$
$=\left.{}_{f}\left\langle 0\right|\left\langle in\right|B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi\left(a^{\dagger}a\right)}\right.\frac{\partial^{m_{1}+n_{1}+m_{2}+n_{2}}}{\partial x_{1}^{m_{1}}\partial y_{1}^{n_{1}}\partial x_{2}^{m_{2}}\partial y_{2}^{n_{2}}}$
$\left.\exp\left[a^{\dagger}x_{1}+b^{\dagger}y_{1}+ax_{2}+by_{2}\right]\right|_{x_{1}=y_{1}=x_{2}=y_{2}=0}$
$\times e^{i\phi\left(a^{\dagger}a\right)}B_{f}B_{1}\left|in\right\rangle\left|0\right\rangle_{f}.$ (B4)
@@
$\begin{array}[]{c}\Gamma_{1}\{\exp\left[\alpha\left(t_{1}+t_{2}+A_{1}+A_{4}\right)\cosh r\right]\\ \times\exp\left[\alpha\left(\tau_{1}+\tau_{2}+A_{2}+A_{3}\right)\sinh r\right]\\ \times\exp[\left(t_{1}\sinh r+A_{2}\cosh r\right)\tau_{1}\cosh r\\ +\left(\tau_{2}\sinh r+A_{1}\cosh r\right)t_{2}\cosh r]\\ \times\exp[\left(t_{1}\cosh r+A_{2}\sinh r\right)\\ \times\left(A_{3}\sinh r+\left(t_{2}+A_{4}\right)\cosh r\right)]\\ \times\exp[\left(\tau_{2}\cosh r+A_{1}\sinh r\right)\\ \times\left(A_{4}\sinh r+\left(\tau_{1}+A_{3}\right)\cosh r\right)]\}\end{array}$ (31i)
$=D_{1}\left(m_{1},n_{1},m_{2},n_{2}\right),$ (B5)
@@
$\Gamma_{1}\left\{\cdot\right\}$
$=\frac{1}{N_{m,n}}\frac{\partial^{2m+2n+m_{1}+n_{1}+m_{2}+n_{2}}}{\partial t_{1}^{m}\partial\tau_{1}^{n}\partial t_{2}^{m}\partial\tau_{2}^{n}\partial x_{1}^{m_{1}}\partial y_{1}^{n_{1}}\partial x_{2}^{m_{2}}\partial y_{2}^{n_{2}}}\left\{\cdot\right\}$
$|_{t_{1}=\tau_{1}=t_{2}=\tau_{2}=x_{1}=y_{1}=x_{2}=y_{2}=0},$ (B6)

### v2

@@ III Phase Sensitivity of Intensity-Difference Detection
Figs. 4(a)–(c) depict linear phase shift sensitivity $\Delta\phi$ as a function of loss rate $l$, compared with SQL and HL. Despite photon loss, phase sensitivity can still surpass SQL, with photon addition improving system robustness. Figs. 5(a)–(c) further illustrate Kerr nonlinear phase shift sensitivity versus loss rate $l$. Comparison with Fig. 4 clearly shows that Kerr nonlinearity enhances phase sensitivity by an order of magnitude, enabling exceedance of the HL threshold even with loss, whereas linear phase shift only surpasses SQL within limited $l$ ranges. However, with fixed $r=0.9$ and $\alpha=1$, phase sensitivity for $m=n=0$ can exceed sub-HL, but not for $m=n=1,2$. This may arise from Kerr nonlinearity sensitivity to other parameters (e.g., phase shift). This situation indicates that without photon addition, phase sensitivity can exceed sub-HL within certain ranges, which is relatively favorable for Kerr nonlinearity.
@@ IV.1 QFI of KMZI in Ideal Case
Quantum Fisher information (QFI) quantifies the maximum information about phase shift $\phi$ that a quantum state can convey after passing through a phase shifter. In our scheme, employing linear and Kerr nonlinear phase shifters within the MZI leads to two corresponding QFI representations, denoted $F_{k}$ ($k=1,2$). QFI is independent of any specific detection scheme and provides an upper bound on classical Fisher information obtainable through $\phi$ estimation. The optimal error bound for phase sensitivity, known as the quantum Cramér–Rao bound (QCRB), is given by 44 ; 45 ; 46 :
@@
$F_{k}=4\bigl[\langle\psi_{\phi}^{\prime}|\psi_{\phi}^{\prime}\rangle-\lvert\langle\psi_{\phi}^{\prime}|\psi_{\phi}\rangle\rvert^{2}\bigr],$ (14)
where $\lvert\psi_{\phi}\rangle=U(\phi,k)B_{1}\lvert\psi\rangle_{\text{in}}$ represents the quantum state before BS2 in the lossless KMZI, and $\lvert\psi_{\phi}^{\prime}\rangle=\partial\lvert\psi_{\phi}\rangle/\partial\phi$. For both linear ($k=1$) and Kerr nonlinear ($k=2$) phase shifts, QFI can be further simplified as 40 :
$F_{k}=4\langle\Delta^{2}n_{a}^{k}\rangle=4\bigl[\langle(n_{a}^{k})^{2}\rangle-\langle n_{a}^{k}\rangle^{2}\bigr],$ (15)
where $\langle\cdot\rangle=\langle\Psi\rvert\cdot\lvert\Psi\rangle$, with $\lvert\Psi\rangle=B_{1}\lvert\text{in}\rangle$. Here, $n_{a}^{k}=(a^{\dagger}a)^{k}$ ($k=1,2$). To facilitate QFI computation, the normal ordering form of $n_{a}^{w}$ is obtained via the operator identity:
$n_{a}^{w}$ $=$ $\frac{\partial^{w}}{\partial x^{w}}\exp\bigl[xa^{\dagger}a\bigr]\Big|_{x=0}$ (16)
$=$ $\frac{\partial^{w}}{\partial x^{w}}\colon\!\exp\bigl[(e^{x}-1)a^{\dagger}a\bigr]\!\colon\Big|_{x=0}.$
@@
$F_{1}=4\bigl[D_{1}(2,0,2,0)+D_{1}(1,0,1,0)-D_{1}(1,0,1,0)^{2}\bigr],$ (17)
@@
$F_{2}$ $=$ $4\bigl[D_{1}(4,0,4,0)+6D_{1}(3,0,3,0)$ (18)
$+7D_{1}(2,0,2,0)+D_{1}(1,0,1,0)$
$-\bigl(D_{1}(2,0,2,0)+D_{1}(1,0,1,0)\bigr)^{2}\bigr].$
@@
As evident from Fig. 6, increasing photon-addition order significantly enhances QFI. Furthermore, comparing Figs. 6(a)–(b) with Figs. 6(c)–(d) clearly shows that the $k=2$ case yields larger QFI values. This demonstrates that both photon-addition operations and Kerr nonlinear phase shifts effectively improve QFI.
@@ IV.2.1 Effects of Photon Loss on QFI in Linear Phase Shift Case ($k=1$)
Practical measurements involve photon losses due to vacuum noise, affecting QFI. For the initial pure state $\lvert\Psi_{S}\rangle$ ($\lvert\Psi_{S}\rangle=B_{1}\lvert\psi\rangle_{\text{in}}$) in a probe system $S$ with a lossy MZI, we introduce environment $E$ orthogonal states $\lvert j_{E}\rangle$ and Kraus operators $\hat{\Pi}_{j}(\phi)$ to characterize $\lvert\Psi_{S}\rangle$ evolution through the phase shifter with photon losses. Fig. 7 illustrates the photon loss within the MZI, providing a visual representation of the model structure under investigation. Following Escher et al.’s methods for calculating QFI in open quantum systems 49 , the quantum states $\lvert\Psi_{S}\rangle$ and environment vacuum noise $\lvert 0_{E}\rangle$, after unitary evolution $U_{S+E}(\phi)$, can be expressed in the extended $S+E$ space:
$\lvert\Psi_{S+E}\rangle$ $=$ $U_{S+E}(\phi)\lvert\Psi_{S}\rangle\lvert 0_{E}\rangle$ (19)
$=$ $\sum_{j=0}^{\infty}\hat{\Pi}_{j}(\phi)\lvert\Psi_{S}\rangle\lvert j_{E}\rangle.$
Although photon losses transform $\lvert\Psi_{S}\rangle$ into a mixed state, this approach allows treating $\lvert\Psi_{S+E}\rangle$ in $S+E$ as a pure state. For the entire purified system, QFI under photon losses can be expressed as:
$F_{L_{1}}\leqslant C_{Q}\bigl[\lvert\Psi_{S}\rangle,\hat{\Pi}_{j}(\phi)\bigr]=4\bigl[\langle\hat{H}_{1}\rangle_{S}-\lvert\langle\hat{H}_{2}\rangle_{S}\rvert^{2}\bigr],$ (20)
where the lower bound of $C_{Q}[\lvert\Psi_{S}\rangle,\hat{\Pi}_{j}(\phi)]$ is shown to be the QFI for the reduced system 49 , and $\hat{H}_{1,2}$ are Hermitian operators defined as:
$\hat{H}_{1}$ $=$ $\sum_{j=0}^{\infty}\frac{d\hat{\Pi}_{j}^{\dagger}(\phi)}{d\phi}\frac{d\hat{\Pi}_{j}(\phi)}{d\phi},$ (21)
$\hat{H}_{2}$ $=$ $i\sum_{j=0}^{\infty}\frac{d\hat{\Pi}_{j}^{\dagger}(\phi)}{d\phi}\hat{\Pi}_{j}(\phi),$ (22)
with $\langle\cdot\rangle_{S}$ denoting expectation value $\langle\Psi_{S}\rvert\cdot\lvert\Psi_{S}\rangle$, representing the average of $\hat{H}_{1,2}$. Thus, $F_{L_{1}}=C_{Q\min}$, and for linear phase shift ($k=1$), the Kraus operator is:
$\hat{\Pi}_{j}(\phi)=\sqrt{\frac{l^{j}}{j!}}e^{i\phi(a^{\dagger}a-\gamma j)}(1-l)^{\frac{a^{\dagger}a}{2}}a^{j},$ (23)
@@
$F_{L_{1}}=\frac{4F_{1}(1-l)\langle n_{a}\rangle}{lF_{1}+4(1-l)\langle n_{a}\rangle}.$ (24)
@@ IV.2.2 Effects of Photon Loss on QFI in Kerr Nonlinear Phase Shift Case ($k=2$)
$\hat{\Pi}_{j}(\phi)=\sqrt{\frac{l^{j}}{j!}}e^{i\phi[(a^{\dagger}a)^{2}-2\mu_{1}a^{\dagger}aj-\mu_{2}j^{2}]}(1-l)^{\frac{a^{\dagger}a}{2}}a^{j},$ (25)
@@
$F_{L_{2}}$ $\leqslant$ $C_{Q}\bigl[\lvert\Psi_{S}\rangle,\hat{\Pi}_{j}(\phi)\bigr]$ (26)
$=$ $4\bigl[K_{1}^{2}\langle\Delta^{2}n_{a}^{2}\rangle_{S}-K_{2}\langle n_{a}^{3}\rangle_{S}$
$+K_{3}\langle n_{a}^{2}\rangle_{S}-K_{4}\langle n_{a}\rangle_{S}$
$-K_{5}\langle n_{a}^{2}\rangle_{S}\langle n_{a}\rangle_{S}-K_{6}\langle n_{a}\rangle_{S}^{2}\bigr],$
@@ V Conclusion
We have proposed and analyzed an enhanced phase estimation scheme combining PA-TMSCS with a KMZI. Our results demonstrate that this approach systematically improves phase sensitivity and QFI, enabling the surpassing of fundamental precision limits. Specifically, the scheme exceeds the SQL and approaches the HL for linear phase shifts, while the Kerr nonlinearity facilitates scaling beyond the SHL. Moreover, the integrated design exhibits enhanced robustness against photon loss, a critical attribute for practical implementations.
This work elucidates the synergistic mechanism between non-Gaussian photon addition and Kerr nonlinearity, which concurrently boosts measurement precision and resilience. The findings establish a valuable theoretical framework for developing practical quantum-enhanced metrology protocols, with potential applications in ultra-precise sensing. Future work may focus on identifying optimal operating parameters and experimental realizations in physical platforms suitable for nonlinear quantum optics.
Acknowledgments
@@
This work is supported by the National Natural Science Foundation of China (Grants No. 12564049 and No. 12104195) and the Jiangxi Provincial Natural Science Foundation (Grants No. 20242BAB26009 and 20232BAB211033), Jiangxi Provincial Key Laboratory of Advanced Electronic Materials and Devices (Grant No. 2024SSY03011), Jiangxi Civil-Military Integration Research Institute (Grant No. 2024JXRH0Y07), as well as the Science and Technology Project of Jiangxi Provincial Department of Education (Grant No. GJJ2404102).
@@
Appendix A Derivation of Transformation Relation for Kerr Nonlinear Phase Shifter Operator
@@
In the scheme proposed in this paper, the Kerr nonlinear phase shifter significantly enhances phase measurement accuracy. The associated calculations can be efficiently performed using the transformation relation of the equivalent operator of the Kerr nonlinear phase shifter, $U(\phi,2)=e^{i\phi(a^{\dagger}a)^{2}}$, as shown in our study. Therefore, this appendix provides a detailed derivation of the transformation relation $U^{\dagger}(\phi,2)aU(\phi,2)=e^{i\phi}e^{i\phi(2a^{\dagger}a)}a$ in Eq. (8). Here, we assume the operator corresponding to the transformation relation of $U(\phi,2)$ is $O$, i.e., $O=U^{\dagger}(\phi,2)aU(\phi,2)$. The matrix of $O$ in the Fock state space can be represented by $O_{n,n^{\prime}}=\langle n\rvert O\lvert n^{\prime}\rangle$, and its corresponding expansion is given by
$O=\sum_{n,n^{\prime}=0}^{\infty}\lvert n\rangle\langle n^{\prime}\rvert O_{n,n^{\prime}}.$ (A1)
@@
By utilizing $\langle n\rvert a^{\dagger}a=\langle n\rvert n$ and $\langle n\rvert a=\langle n+1\rvert\sqrt{n+1}$, the matrix element $O_{n,n^{\prime}}$ in the aforementioned equation can be further calculated as
$O_{n,n^{\prime}}=\sqrt{n+1}e^{i\phi(2n+1)}\delta_{n+1,n^{\prime}},$ (A2)
and by substituting it into Eq. (A1), the transformation relation of $U(\phi,2)$ can be deduced as follows
$O$ $=e^{i\phi}\sum_{n=0}^{\infty}\lvert n\rangle\langle n+1\rvert\sqrt{n+1}e^{i2n\phi}$
$=e^{i\phi}\sum_{n=0}^{\infty}\lvert n\rangle\langle n\rvert e^{i2n\phi}a=e^{i\phi}e^{i\phi(2a^{\dagger}a)}a,$ (A3)
@@
where we have used the completeness of the Fock state $\sum_{n=0}^{\infty}\lvert n\rangle\langle n\rvert=1$.
@@
Appendix B Specific Derivations and Expressions for $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ under $k=1$ Condition
@@
In order to obtain the mean values $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ in Eq. (10) in a simple way, the general formula is introduced, i.e.,
${}_{f}\langle 0\rvert\langle\text{in}\rvert B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi a^{\dagger}a}$
$\times\bigl(a^{\dagger m_{1}}b^{\dagger n_{1}}a^{m_{2}}b^{n_{2}}\bigr)e^{i\phi a^{\dagger}a}B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}$
$=D_{1}(m_{1},n_{1},m_{2},n_{2}).$ (B1)
@@
The quantum state used in the above formula to calculate the average value is the state after passing through the phase shifter, and the output state still needs to pass through BS2. Thus using the transformation relation Eqs. (6) and (B1) one can obtain $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ respectively as
$\langle I_{D}\rangle=i\bigl[D_{1}(1,0,0,1)-D_{1}(0,1,1,0)\bigr],$ (B2)
@@
$\langle I_{D}^{2}\rangle$ $=D_{1}(0,1,0,1)+D_{1}(1,0,1,0)+2D_{1}(1,1,1,1)$
$\quad-D_{1}(2,0,0,2)-D_{1}(0,2,2,0).$ (B3)
@@
$D_{1}(m_{1},n_{1},m_{2},n_{2})$
$={}_{f}\langle 0\rvert\langle\text{in}\rvert B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi a^{\dagger}a}\frac{\partial^{m_{1}+n_{1}+m_{2}+n_{2}}}{\partial x_{1}^{m_{1}}\partial y_{1}^{n_{1}}\partial x_{2}^{m_{2}}\partial y_{2}^{n_{2}}}$
$\quad\exp\bigl[a^{\dagger}x_{1}+b^{\dagger}y_{1}+ax_{2}+by_{2}\bigr]\Big|_{x_{1}=y_{1}=x_{2}=y_{2}=0}$
$\quad\times e^{i\phi a^{\dagger}a}B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}.$ (B4)
@@
$\Gamma_{1}\Bigl\{\exp\bigl[\alpha(t_{1}+t_{2}+A_{1}+A_{4})\cosh r\bigr]$
$\times\exp\bigl[\alpha(\tau_{1}+\tau_{2}+A_{2}+A_{3})\sinh r\bigr]$
$\times\exp\bigl[(t_{1}\sinh r+A_{2}\cosh r)\tau_{1}\cosh r$
$\quad+(\tau_{2}\sinh r+A_{1}\cosh r)t_{2}\cosh r\bigr]$
$\times\exp\bigl[(t_{1}\cosh r+A_{2}\sinh r)$
$\quad\times(A_{3}\sinh r+(t_{2}+A_{4})\cosh r)\bigr]$
$\times\exp\bigl[(\tau_{2}\cosh r+A_{1}\sinh r)$
$\quad\times(A_{4}\sinh r+(\tau_{1}+A_{3})\cosh r)\bigr]\Bigr\}$
$=D_{1}(m_{1},n_{1},m_{2},n_{2}),$ (B5)
@@
$\Gamma_{1}\{\cdot\}$ $=\frac{1}{N_{m,n}}\frac{\partial^{2m+2n+m_{1}+n_{1}+m_{2}+n_{2}}}{\partial t_{1}^{m}\partial\tau_{1}^{n}\partial t_{2}^{m}\partial\tau_{2}^{n}\partial x_{1}^{m_{1}}\partial y_{1}^{n_{1}}\partial x_{2}^{m_{2}}\partial y_{2}^{n_{2}}}\{\cdot\}$
$\quad\Big|_{t_{1}=\tau_{1}=t_{2}=\tau_{2}=x_{1}=y_{1}=x_{2}=y_{2}=0},$ (B6)